From acd@roe.ac.uk Tue Sep 9 14:01:41 2014 Date: Tue, 9 Sep 2014 11:32:48 +0000 From: Clive Davenhall To: Bob Mann , Andy Lawrence , "Read, Mike" , Eckhard Sutorius , Mark S. Holliman , Nicholas Cross , Nigel C. Hambly , Rob Blake , Ross Collins , Stelios Voutsinas , Dave Morris , Clive Davenhall Cc: "dante@astro.puc.cl" , James Dunlop , "j.p.emerson@qmul.ac.uk" , ktn keithnoddle.org , "Lorenzo.Rimoldini@unige.ch" , "mcioni@aip.de" , "matt.jarvis@astro.ox.ac.uk" , "mike@ast.cam.ac.uk" , "nigel.metcalfe@durham.ac.uk" , "norman@astro.gla.ac.uk" , "pwl@star.herts.ac.uk" , "rgm@ast.cam.ac.uk" , "Simon.Dye@nottingham.ac.uk" , "s.j.warren@ic.ac.uk" , "tkerr@jach.hawaii.edu" , "tom.shanks@durham.ac.uk" Subject: WFAU VDFS Science Archives progress meeting minutes WFAU VDFS Science Archives progress meeting minutes, 9 September 2014 Present: MSH, NJC, ETWS, RPB, RSC, ACD Apologies: RGM, MAR, NCH, KTN, AL, STV DoNM: 10am Friday 19 September 2014 in the Villas meeting room, MAR in the chair (room is booked). Actions discharged this time: ---------------------------- ACTION: RPB to put all prices of additional items associated with the backup policy on the Twiki pages. Done (some weeks ago). ACTION: MSH to produce a costed list of the additional kit necessary to permit wholesale reprocessing of UKIDSS/VISTA. Done. ACTION: RPB to make the necessary Spitzer SMC/VMC neighbour table entries. Done. Actions partly discharged but continuing: ---------------------------------------- ACTION: RPB (was MAR) to have a look at the status of non-survey releases for registered projects. MAR had compiled a list of non-survey projects that could be released and sent it to RPB. RPB had yet to make these data available. Actions carried forward from 8/8/2014 meeting: ---------------------------------------------- ACTION: RGM to investigate the size and content (how does it differ from DR9?) of DR10 to see if it's worth us hosting and whether it is necessary to keep all of the other releases. Continues. RGM had started a wiki page for WFAU SDSS strategy http://apache.roe.ac.uk/twiki/bin/view/WFAU/SDSSDRPlan. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues: this one still has low priority. ACTION: ACD to review VVV brainstorming Twiki page and decide who to chase about progressing their assigned actions. Continues. ACTION: RPB to write a script to backup the SVN repository. Continues. ACTION: ACD, RGM: review finances. Continues. Specific points and new actions: ------------------------------- Project management: * Nick Rowell would be joining WFAU to work on the GDFS on 1/10/14. * Office space was tight. NR would be moving into R3 and consequently SV would be moving into R12 or R13. PMW had also asked about temporarily having desk space in R12 or R13. * The Open Day is scheduled for the end of the month (27-28/9/14). WFCAM, VISTA and VST updates: * Nothing to report. Comments and issues arising from CASU minutes: * No minutes had been received since the previous meeting. Review of Deadlines: * Deadlines listed for the past couple of months were reviewed: August 15: VMC P92 team release The VMC P92 team release occurred on 4/9/14. 29: UKIDSS GPS release into DR10 CU19 had now been started. The revised release date is now around 15/9/14. 29: GES iDR3 [data] likely to be available [from Cambridge] The iDR3 data had not yet been received but were expected soon. September 5: VVV DR2 ESO release The process to convert the WFAU VVV DR2 release into the format required by ESO is currently running. 5: VIDEO Team release 2 CASU are still reprocessing the data; they are believed to be well-advanced with this task. Currently it is not possible to give an estimate of the release date. 15: VIDEO ESO release Currently this release is delayed. 19: Aim to have started CU6 on VVV DR3 so that it is complete in time to run CU19 over Christmas. Likely to be delayed. Indeed, it may now be impossible to have VVV DR3 available before the Vatican VVV meeting in May 2015. Networking: * The networks for WFAU operations were working normally. * Network connections in office R3 were briefly discussed. A number of spare ports had been identified. Some of these were IP telephone ports but it was believed that these could be reconfigured as computer ports. RSC was happy to liaise with IT Support over this issue. WSA/VSA/OSA/GES Operations: * ETWS reported that: - The objID update for the VVV is running smoothly along. If it scales well the last batch of updates should finish in less than 10 days time. - OSA catalogue ingests are running well. Unfortunately the shares on Sahure and Userkaf filled up faster than the ingests could be run, so the CU4 runs broke down. The runs on the nights that were not processed have been restarted. - Transferred, JPGed, and ingested the metadata for the latest 4 months of WFCAM data released to us (April-July 2014). * ACD reported on GES: - GES iDR2 was now largely complete. - The remaining task for iDR2 was the nightly radial velocities, which he hoped to look at next week. - The documentation (Web pages) had been updated for iDR2 and should require little change for iDR3. - iDR3 data were expected from Cambridge soon. - ACD had registered to attend the GES conference in Porto, Portugal in November, where he hoped to give a talk and arrange a hands-on tutorial session. Hardware and Systems: * MSH reported that: - The network was slowly being upgraded. - There were a few machines for which the existing network connection must be removed before upgrading to the new one. So, down-time will be required and would be arranged to minimise disruption. - The RAID array on Kakai was running slowly and would be rebuilt next week. Software: * NJC noted that: - He had resolved the last issue with the VVV DR2 ESO Release; the software is now running (as noted above). Once it finishes some checking will be needed. - The software for the next VIDEO team release has been checked and some inconsistencies found between the OB frames that CASU supplied and those supplied to the VIDEO team, which is not a big surprise because CASU were known to be re-sky-subtracting the VIDEO images, though it was assumed that this had been done for the v1.3 release. - Some minor software changes had been made for the new VIDEO mosaics, mainly to do with setting up for the new disks to host them. - Finally there had been changes to the programme setup software to deal with the new list-driven u-band data coming into the OSA. Some of the changes reordered the filter list in WSA/VSA source tables, which could have caused all sorts of problems, so changes were made to deal with this and also to work with the OSA * ETWS reported some minor fixes on the OSA side of processing code. Survey Data Release: * VMC P92 had been released on 4/9/14. Non-survey Data Release: * None. Astrogrid deployment, VO & Data Analysis services: * The deployed systems remained operational; nothing to report. Miscellaneous: * ACD reminded the meeting about the forthcoming 'The Universe of Digital Sky Surveys' Sky Survey conference to be held in Naples during November. This meeting is highly relevant to WFAU's work. Unfortunately NJC was unable to go (he had some feeble excuse involving his wife's graduation). RSC noted that the conference clashed with a week of EUCLID meetings, making it difficult for people working on EUCLID to attend. ACD was reluctant to go because he hoped to go to the GES meeting in November. He noted that the call for volunteers was still open and that if none were forthcoming by early next week he would organise a press-gang. ACD, Original: 5/9/13, Revised: 5/9/14. -- Scanned by iCritical.